How many millilitres of 0.5 M H2SO4 are needed to dissolve 0.5 g of copper (II) carbonate?

Grade:upto college level

4 Answers

Aditi Chauhan
9 years ago
Sol . The balanced chemical reaction is : CuCO3 + H2SO4 → CuSO4 + H2O + CO2 Millimol of CuCO3 = (0.5 × 1000)/123.5 = 4.048 ⇒ millimol of H2SO4 required = 4.048 ∵ millimol = Molarity × Volume (in mL) ⇒ Vol = 4.048/0.50 = 8.1 mL
mayank
6 years ago
Sol . The balanced chemical reaction is : CuCO3 + H2SO4 → CuSO4 + H2O + CO2 Millimol of CuCO3 = (0.5 × 1000)/123.5 = 4.048 ⇒ millimol of H2SO4 required = 4.048 ∵ millimol = Molarity × Volume (in mL) ⇒ Vol = 4.048/0.50 = 8.1 mL
Ashwin Sheoran
5 years ago
 balanced chemical reaction is : 
CuCO3 + H2SO4 → CuSO4 + H2O + CO2 
Millimol of CuCO3 = (0.5 × 1000)/123.5 
= 4.048 
⇒ millimol of H2SO4 required = 4.048 
∵ millimol = Molarity × Volume (in mL) 
⇒ Vol = 4.048/0.50 = 8.1 mL 
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

The balanced chemical reaction is :
CuCO3 + H2SO4 → CuSO4 + H2O + CO2
Millimol of CuCO3 = (0.5 × 1000)/123.5
= 4.048
⇒ millimol of H2SO4 required
= 4.048
∵ millimol = Molarity × Volume (in mL)
⇒ Vol = 4.048/0.50
= 8.1 mL

Thanks and Regards